A body of mass 8.0 kg is traveling at 2.0 m/s under the influence of no external force. At a certain instant an internal explosion occurs, splitting the body into two chunks of 4.0 kg mass each; 16 J of translational kinetic energy are imparted to the two-chunk system by the explosion. Neither chunk leaves the line of the original motion. Determine the speed and direction of motion of each of the chunks after the explosion.

Grade:upto college level

1 Answers

Hasan Naqvi
8 years ago
Original kinetic energy= 1/2mv2 =  16J
Energy imparted due to explosion = 16J
Total energy = 32J
 
½ m1v12 + 1/2 m2v22  = 32J
2 v12 + 2v2= 32J
v12 + v2= 16J
 
Also m1v1 + m2v= mv (conservation of linear momentum)
4v1 + 4v2 = 8 x 2
v+ v2 = 4
(v+ v2 )2= 16
i.e. v1v2 = 0
 
So one of the chunk stops at rest, while the other chunk moves with a velocity 4m/s in the same direction.